OpenStudy (sarahann):

Find the surface area of a cylinder with diameter 10 and height 4.

OpenStudy (cwrw238):

surface area of the 2 ends = 2*pi*r^2 surface area of the curved part = 4 * 2*pi*r r = 1/2 * diameter = 5